24,In June, employment was 9,553,800. In July it was 9,574,800. What was thelmonthly percet increase in employment at an annual rate? a. 0.22% b. 2.67% c. 2.64% 4
Answer #1

b. 2.67%

Firstly, we find the monthly increase.
Monthly increase = (employment in July - employment in June)/employment in June
= (9,574,800-9,553,800)/9,553,800 = 21,000/9,553,800 = 0.0022

Annual growth rate = [(1+r)^n] - 1
where r = monthly increase = 0.0022
n = total month in a year = 12
So, Annual growth rate = [(1+0.0022)^12] - 1 = Annual growth rate = [(1.0022)^12] - 1 = 1.0267 - 1 = 0.0267
Converting this to percentage we get 0.0267*100 = 2.67%